In a World Atlas study, 10% of people have blue eye color. Lane decided to observe 35 people and she concluded 5 people had blue eyes. Calculate the z-score.

1. 0.1428

2. 0.8452

3. 0.0041

4. 0.0430

4. 0.0430
Answer #1

Answer: 2. 0.8452

Solution:

The required formula is given as below:

Z = (p̂ - p)/sqrt(pq/n)

We are given

x = 5

n = 35

p̂ = x/n = 5/35 = 0.142857

p = 0.10

q = 1 - p = 1 - 0.10 = 0.90

Z = (0.142857 - 0.10)/sqrt(0.10*0.90/35)

Z = 0.845151

Z = 0.8452